# Identifiable learning of dissipative dynamics

### v1

@@ 4.1 Linear Case
We begin by validating the effectiveness of I-OnsagerNet in a simple, analytically tractable linear diffusion process. Consider the 2D linear SDE [47]:
@@
where the (constant) matrices $M$ and $S$ are symmetric positive-definite and $W$ is antisymmetric with $W=\lambda W_{0}$, where $W_{0}$ is the standard antisymmetric matrix in $2$ dimensions. Linearity yields the closed-form stationary density $\rho=\mathcal{N}(0,S^{-1})$. From (6) we compute the global EPR [19]:
@@
This SDE coincides exactly with I-OnsagerNet in the linear regime, with the irreversible component $W\,S\,\mathbf{Z}$. While I-OnsagerNet employs non-linear networks and is a non-linear model, the analytical tractability of this linear system provides exact expressions for the stationary density, the irreversible component, and the EPR, offering a rigorous ground truth for quantitative benchmarking.
Figure 2: Results for linear system. (a) The learned potential and the negative log of the stationary density, each shifted by subtracting their value at input zero, ensuring that both functions attain zero at the origin. (b) First component for the time-irreversible drift of learned and exact system for $\lambda=1$. (c) Comparison of the global EPR calculated via four approaches. The exact value is computed using (8); Est.1 is calculated by applying a state-space discretization to approximate the definition in (5); Est.2 and learned $\dot{S}_{\text{tot}}$ are obtained via Monte Carlo approximation of (6), using the exact and learned equations, respectively.
We generate trajectory data by simulating this linear SDE and use these data to train I-OnsagerNet. The results presented in Fig. 2 provide a direct validation of our approach. This confirms that our method faithfully reconstructs both the stationary density and the irreversible component of the dynamics. In addition, the global EPR across a range of values of $\lambda$, the parameter that controls the amount of irreversibility, shows excellent agreement with theoretical predictions. Computing irreversibility from discrete trajectory observations presents significant challenges due to the difficulty in estimating path measures. One common approach involves discretizing the state space (the Appendix). However, combined errors arising from both temporal and spatial discretization impair the reliability of the resulting statistical estimates, and the computational cost increases exponentially with the system dimension. As shown in Fig. 2 (c), the direct computation of the EPR (Est.11) captures only the qualitative trends of the exact EPR without matching its quantitative values.
Figure 3: Simulation setup, variations across random seeds and barrier heights of the learned potential energy for polymer stretching dynamics, (a) Linear bead-rod polymer chains are simulated using a Brownian dynamics framework under a planar elongational flow, whose strength is characterized by the strain rate $\dot{\varepsilon}$. (b) Relative variation ($\times 10^{3}$) in learned potential forces over 44 random seeds at different strain rate $\dot{\varepsilon}$. (c-h) Projections of the learned potential energy at $\dot{\varepsilon}=4.63$ onto the $Z_{1}\text{–}Z_{2}$ (c,f), $Z_{1}\text{–}Z_{3}$ (d,g), and $Z_{2}\text{–}Z_{3}$ (e,h) planes, comparing I-OnsagerNet (c-e) and OnsagerNet (f-h) across 33 seeds. (i.j) Energy landscape barrier heights of I-OnsagerNet and OnsagerNet averaged over $4$ independent experiments; error bars indicate one standard deviation.
Figure 4: EPR for polymer stretching dynamics. (a) The global EPR obtained via I-OnsagerNet for varying strain rate $\dot{\varepsilon}$, averaged over $4$ independent experiments with error bars showing one standard deviation. (b) Illustration of the experimental validation and coarse-graining of the global EPR based on device-level observations. (c-f) The total EPR $\dot{s}_{\text{tot}}$ on test trajectories projected onto the $Z_{1}\text{–}Z_{2}$ and $Z_{1}\text{–}Z_{3}$ planes. The background is the corresponding potential energy surface. One example trajectory from the test dataset are overlaid on the $Z_{1}\text{–}Z_{3}$ panels.
@@ 4.2 Polymer Stretching Dynamics
We next demonstrate our approach by modeling the temporal evolution of polymer chain extension under elongational flow. Owing to its ubiquity in polymer processing, this physical process is fundamental to the study of polymer rheology and has been a subject of extensive research within the polymer physics community [4, 20, 24]. Landmark experiments and simulations at the single-molecule level have revealed a profound heterogeneity, in which identical polymer chains exhibit starkly different stretching dynamics, and the statistical nature of this heterogeneity is acutely sensitive to system parameters such as flow strength [62, 67, 44]. A recent deep learning strategy seeks to capture these dynamics by constructing macroscopic descriptors and their effective energy landscape [13]. However, the complexity and inherent stochasticity of polymer dynamics lead to identifiability issues in these models. The learned energy function lacks a unique physical definition, leading to inconsistent solutions even for identical datasets; thus the inferred non-equilibrium components fail to quantitatively capture the system’s entropy production. This ambiguity has made it difficult to systematically investigate the influence of physical parameters on polymer dynamics. In this section, we show that our I-OnsagerNet is able to overcome these limitations, successfully capturing heterogeneous dynamics across various flow strengths while providing a physically meaningful energy landscape and a quantitative measure of entropy production.
@@
We simulate the stretching of a single polymer chain in a planar elongational flow for various flow strengths, characterized by the strain rate $\dot{\varepsilon}$ (Fig. 3 (a)) . The chain consists of 300 coarse-grained beads connected by rigid rods, resulting in a system with 900 degrees of freedom before imposing constraints to neglect inertial effects. The polymer extension, defined as the projected length along the elongational axis, provides a macroscopic view of the microscopic heterogeneity of the polymer population and is therefore chosen as our coordinate of interest $Z_{1}$. Our approach constructs two additional closure coordinates alongside $Z_{1}$ following previous approaches to model reduction [13], thereby establishing a three-variable I-OnsagerNet for each value of the strain rate. Validated on various initial chain conformations, the trained models demonstrate precise prediction of the chain extension evolution across all cases (the Appendix).
@@
Figure 3 provides a numerical validation of the identifiability of the proposed I-OnsagerNet. Specifically, Panel (b) demonstrates that the relative variation in the learned potential forces across different random seeds is significantly smaller for our method compared to standard OnsagerNet. Panels (c-h) further illustrate this by showing that I-OnsagerNet consistently produces identical projected potential surfaces at $\dot{\varepsilon}=4.63$, irrespective of random initialization. Here the projections are obtained via minimization (e.g., $V(Z_{1},Z_{2})=\min_{Z_{3}}V(Z_{1},Z_{2},Z_{3})$), which closely approximates the marginal potential at low temperatures.
@@
Having established the identifiability, we now investigate the energy landscape of the polymer stretching dynamics under a range of strain rates $\dot{\varepsilon}$. The barrier height provides a quantitative measure of the energetic difficulty associated with transitions between metastable configurations [33, 71]. In the context of polymer stretching, it represents the effective free-energy barrier associated with the flipping of chain ends against the imposed elongational flow. Although it does not exactly correspond to the transition rates in our nonequilibrium setting, it serves as a qualitative descriptor of the effective energy landscape. Since I-OnsagerNet directly provides the logarithm of the stationary density, it allows for a straightforward computation of the barrier height (the Appendix). As shown in Fig. 3(i), the barrier height associated with polymer stretching exhibits a super-linear increase with the strain rate $\dot{\varepsilon}$. In contrast, the OnsagerNet approach fails to produce consistent energy functions for fixed parameters, resulting in barrier heights that show no clear or systematic dependence on the strain rate, as illustrated in Fig. 3(j).
@@
Next, we examine the irreversibility by computing the global EPR, $\dot{S}_{\text{tot}}$. The log-log plot in Fig. 4(a) reveals a clear scaling relationship with the strain rate $\dot{\varepsilon}$. Specifically, $\dot{S}_{\text{tot}}$ exhibits sub-linear growth, tending towards a plateau in both the weak- and strong-driving regimes. The fact that $\dot{S}_{\text{tot}}$ is consistently positive for all strain rates provides evidence that the polymer system operates far from equilibrium.
@@
To understand the origin of this global irreversibility, we subsequently measure the total EPR. Fig. 4 (c-f) shows distinct dynamical regimes governed by the strain rate $\dot{\varepsilon}$ (see the Appendix for results at other $\dot{\varepsilon}$ values). At lower magnitudes of $\dot{\varepsilon}$, the total EPR exhibits minimal temporal variation along the individual trajectories of the polymers. In contrast, for larger values of $\dot{\varepsilon}$, the local EPR displays pronounced spatiotemporal heterogeneity during the unfolding process. Specifically, the total EPR is significantly suppressed at both the initial state (folded polymer) and the final state (fully extended conformation), yet it peaks dramatically during the mid-unfolding phase. This strain rate-dependent bimodal distribution of total EPR highlights the role of non-equilibrium dynamics in driven polymer unfolding, where maximum irreversibility occurs during mid-trajectory unfolding events.
@@
Finally, we conduct single-molecule real-world experiments to verify the qualitative results that the polymer system operates far from equilibrium. A property of the global EPR is that the EPR computed from projected dynamics is a lower bound for that of the full system (the Appendix). Leveraging this property, we can assess non-equilibrium behavior directly from device-level observations by evaluating the EPR of the measured coordinates. We track the stretching dynamics of individual DNA molecules subjected to a planar elongational field, where the strain rate is regulated by the applied voltages (the Appendix). Configurations in which the DNA chain is fully extended are selected to ensure that the system had reached a steady state. From the image-recorded configurations, we extract the time series of the horizontal coordinates of the left and right endpoints, corresponding to the atomic positions at both termini of the DNA molecule. Using only these two-dimensional trajectory data, we numerically compute the global EPR based on the definition (5). This procedure is illustrated in Fig. 4 (b). The resulting EPR values, $71.41$ at $60V$ and $68.99$ at $30V$, remain strictly positive even with partial information, consistent with the non-equilibrium behavior predicted by I-OnsagerNet. These findings highlight the validity of our approach in extracting physically meaningful insights from real polymer dynamics.
@@ 4.3 Stochastic Gradient Langevin Dynamics
Stochastic gradient langevin dynamics (SGLD) [75] is a cornerstone algorithm in modern machine learning, bridging stochastic optimization and Bayesian sampling. It extends standard mini-batch gradient descent by injecting isotropic Gaussian noise to sample from the target distribution:
@@
While the injected noise is designed to drive the system toward an equilibrium Gibbs-Boltzmann distribution, the stochasticity inherent in mini-batch gradient approximation complicates this picture. In the continuous-time limit, the mini-batching introduces a state-dependent covariance into the diffusion term [48]:
@@
We now apply our method to analyze SGLD in the context of least squares regression (linear) and independent components analysis (nonlinear) (the Appendix) to measure the irreversibility and validate the connection between irreversibility and sampling bias. Fig. 5 quantifies the impact of mini-batch size on SGLD irreversibility, measured by the global EPR, across varying batch sizes. Across two distinct systems, the global EPR consistently decreases with increasing mini-batch size at small scales, and plateaus at larger scales due to numerical precision constraints. These results indicate that increasing mini-batch size progressively suppresses non-equilibrium dynamics in SGLD. At sufficiently large batch sizes, negligible global total EPR values indicate that the SGLD performs near detailed balance, and that the non-equilibrium bias vanishes.
@@
Furthermore, we compute the MMD between the invariant distributions from different batch sizes and that of full-batch SGLD. As shown in Fig. 5 (a, c), the MMD trend as a function of batch size is highly consistent with that of the EPR (see the Appendix for theoretical justification). For comparison, we test a heuristic convergence metric based on the MMD between distributions from consecutive mini-batch sizes (i.e., batch sizes $b$ and $b/2$). This baseline metric can be misleading, showing trends inconsistent with actual convergence or lagging as an indicator, as shown in Fig. 5 (b, d). These findings demonstrate that I-OnsagerNet constructed using only sampling trajectories can effectively compute the EPR as a physically meaningful diagnostic of sampling bias.
5 Discussion and Outlook
@@
This work introduces I-OnsagerNet, a framework that unifies physical interpretability with rigorous identifiability for complex dissipative dynamics. By extending the Onsager principle to stochastic and far-from-equilibrium systems, it ensures a unique energy landscape and a clear decomposition of the drift force into reversible and irreversible components. This structure enables key thermodynamic quantities, particularly the entropy production rate that characterizes time-irreversibility, to be directly inferred from data. Applications to polymer stretching and stochastic gradient Langevin dynamics demonstrate the versatility of I-OnsagerNet as a general tool for uncovering unknown dynamics in non-equilibrium phenomena. With its identifiable and physically consistent representation of dynamics, the framework opens new avenues for data-driven discovery and control of complex systems, offering a unified perspective on diverse nonequilibrium processes across scientific disciplines [23, 76, 79]. Building on these foundations, several promising directions remain for further exploration. A natural extension is to integrate operator learning architectures [54, 49, 43], enabling the treatment of systems with varying parameters and dynamic boundary conditions. Another direction is to design equilibrium-preserving sampler and numerical difference scheme by systematically quantifying and reducing irreversibility. Finally, our current framework inherently focuses on macroscopic dynamics learned from coarse-grained descriptions. Extending it to learn directly from mesoscopic or microscopic information through an operator-level formulation offers an exciting pathway toward broader and more versatile applications.
6 Methods
In this section, we detail the architecture of the I-OnsagerNet modules and outline the practical implementation for learning.
6.1 Neural Network Parameterization
$W(\mathbf{Z})=\sum_{d=1}^{D-1}H_{d}(\mathbf{Z})J_{d},$
where $U:\mathbb{R}^{D}\rightarrow\mathbb{R}^{m}$ is a neural network, $\gamma_{ij}$ are trainable parameters, and $\beta_{1},\beta_{2},\beta_{3}>0$ are fixed scaling factors. This architecture ensures that the potential function satisfies both the integrability condition and the appropriate growth requirements specified in the Appendix. The expressivity is achieved through the embedded neural network $U$ within this constrained functional form. In our implementation, we set $m=32$ and use a network $U$ with one hidden layer of 128 neurons. We employ the shifted ReQU activation in polymer dynamics and tanh in the SGLD context.
6.2 Learning Algorithm
Each trajectory is assumed to be a realization of an underlying SDE starting from the initial condition $\mathbf{Z}^{(n)}_{0}$. Our objective is to learn the functions $V$, $W$, and $\sigma$ that govern the dynamics of the I-OnsagerNet within this deep learning framework.
As described in the previous section, we model $V_{\theta}$, $W_{\theta}$, and $\sigma_{\theta}$ with neural networks, resulting in the following parametric SDE:
where
We will use the Maximum Likelihood Estimation (MLE) to learn the parameter $\theta$ [61, Sec. 5.3]. Let $p(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$ be the transition density of the process (i.e., the conditional probability density of the system being in state $\mathbf{Z}$ at time $\Delta t$, given it started at $\mathbf{Z}_{0}$ at time 0). The MLE objective is:
where $\Delta t_{m}=t_{m}-t_{m-1}$. Since the exact transition density is typically intractable for general nonlinear diffusion processes, we approximate it using a numerical approximation $p_{h}(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$ based on the method developed in [80]. The practical optimization problem thus becomes:
Specifically, the one-step scheme from [80] leads to a Gaussian approximation for the transition density:
$p_{h}(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$
This approximation, corresponding to the Euler-Maruyama scheme, has been widely used in related works [13, 22]. We employ the Adam optimizer [42] for training, with detailed hyperparameter settings provided in the the Appendix.
Appendix
Appendix A Theoretical Results

### v2

@@ 2.2 Linear Case
We begin by validating the effectiveness of UIDD in a simple, analytically tractable linear diffusion process. Consider the 2D linear SDE [LelievreNierPavliotis2013]:
@@
where the (constant) matrices $M$ and $S$ are symmetric positive-definite and $W$ is antisymmetric with $W=\lambda W_{0}$, where $W_{0}$ is the standard antisymmetric matrix in $2$ dimensions. Linearity yields the closed-form stationary density $\rho=\mathcal{N}(0,S^{-1})$. This SDE coincides exactly with UIDD in the linear regime, with the irreversible component $W\,S\,\mathbf{Z}$. From Eq. (5) we compute the global EPR [da2023entropy]:
@@
While UIDD employs non-linear networks and is a non-linear model, the analytical tractability of this linear system provides exact expressions for the stationary density, the irreversible component, and the EPR, offering a rigorous ground truth for quantitative benchmarking.
Figure 2: Results for linear system. (a) The learned potential and the negative log of the stationary density, each shifted by subtracting their value at input zero, ensuring that both functions attain zero at the origin. (b) First component for the time-irreversible drift of learned and exact system for $\lambda=1$. (c) Comparison of the global EPR calculated via four approaches. The exact value is computed using Eq. (7); Est.1 is calculated by applying a state-space discretization to approximate the definition in Eq. (4); Est.2 and learned $\dot{S}_{\text{tot}}$ are obtained via Monte Carlo approximation of Eq. (5), using the exact and learned equations, respectively.
We generate trajectory data by simulating this linear SDE and use these data to train UIDD. The results presented in Fig. 2 provide a direct validation of our approach. This confirms that our method faithfully reconstructs both the stationary density and the irreversible component of the dynamics. In addition, the global EPR across a range of values of $\lambda$, the parameter that controls the amount of irreversibility, shows excellent agreement with theoretical predictions. Computing irreversibility from discrete trajectory observations presents significant challenges due to the difficulty in estimating path measures. One common approach involves discretizing the path space (Methods D.1). However, combined errors arising from both temporal and spatial discretization impair the reliability of the resulting statistical estimates, and the computational cost increases exponentially with the system dimension. As shown in Fig. 2 (c), the computation of the EPR via definition (Est.11) captures only the qualitative trends of the exact EPR without matching its quantitative values.
@@ 2.3 Polymer Stretching Dynamics
We next demonstrate our approach by modeling the temporal evolution of polymer chain extension under elongational flow. Owing to its ubiquity in polymer processing, this physical process is fundamental to the study of polymer rheology and has been a subject of extensive research within the polymer physics community [bird1987dynamics, de1974coil, doi1988theory]. Landmark experiments and simulations at the single-molecule level have revealed a profound heterogeneity, in which identical polymer chains exhibit starkly different stretching dynamics, and the statistical nature of this heterogeneity is acutely sensitive to system parameters such as flow strength [larson2005rheology, perkins1997single, smith1998response]. A recent deep learning strategy seeks to capture these dynamics by constructing macroscopic descriptors and their effective energy landscape [chen2024constructing]. However, the complexity and inherent stochasticity of polymer dynamics lead to identifiability issues in these models. The learned energy function lacks a unique physical definition, leading to inconsistent solutions even for identical datasets; thus the inferred non-equilibrium components fail to quantitatively capture the system’s entropy production. This ambiguity has made it difficult to systematically investigate the influence of physical parameters on polymer dynamics. In this subsection, we show that our UIDD is able to overcome these limitations, successfully capturing heterogeneous dynamics across various flow strengths while providing a physically meaningful energy landscape and a quantitative measure of entropy production. This capability further reveals scaling behaviors that were unattainable in existing approaches [chen2024constructing] due to the non-identifiability of potential functions across different flow strengths.
@@
We simulate the stretching of a single polymer chain in a planar elongational flow for various flow strengths, characterized by the strain rate $\dot{\varepsilon}$ (Fig. 3 (a)) . The chain consists of 300 coarse-grained beads connected by rigid rods, resulting in a system with 900 degrees of freedom before imposing constraints to neglect inertial effects. The polymer extension, defined as the projected length along the elongational axis, provides a macroscopic view of the microscopic heterogeneity of the polymer population and is therefore chosen as our coordinate of interest $Z_{1}$. Our approach constructs two additional closure coordinates alongside $Z_{1}$ following previous approaches to model reduction [chen2024constructing], thereby establishing a three-variable UIDD for each value of the strain rate. Validated on various initial chain conformations, the trained models demonstrate precise prediction of the chain extension evolution across all cases (Methods D.4).
Figure 3: Simulation setup, variations across random seeds and barrier heights of the learned potential energy for polymer stretching dynamics, (a) Linear bead-rod polymer chains are simulated using a Brownian dynamics framework under a planar elongational flow, whose strength is characterized by the strain rate $\dot{\varepsilon}$. (b) Relative variation ($\times 10^{3}$) in learned potential forces over 44 random seeds at different strain rate $\dot{\varepsilon}$. (c-h) Projections of the learned potential energy at $\dot{\varepsilon}=4.63$ onto the $Z_{1}\text{-}Z_{2}$ (c,f), $Z_{1}\text{-}Z_{3}$ (d,g), and $Z_{2}\text{-}Z_{3}$ (e,h) planes, comparing UIDD (c-e) and OnsagerNet (f-h) across 33 seeds. Here the projections are obtained via minimization (e.g., $V(Z_{1},Z_{2})=\min_{Z_{3}}V(Z_{1},Z_{2},Z_{3})$), which closely approximates the marginal potential at low temperatures. (i.j) Energy landscape barrier heights of UIDD and OnsagerNet averaged over $4$ independent experiments; error bars indicate one standard deviation.
Figure 4: EPR for polymer stretching dynamics. (a) The global EPR obtained via UIDD for varying strain rate $\dot{\varepsilon}$, averaged over $4$ independent experiments with error bars showing one standard deviation. (b) Illustration of the experimental validation and coarse-graining of the global EPR based on device-level observations. (c-f) The total EPR $\dot{s}_{\text{tot}}$ on test trajectories projected onto the $Z_{1}\text{-}Z_{2}$ and $Z_{1}\text{-}Z_{3}$ planes. The background is the corresponding potential energy surface. One example trajectory from the test dataset are overlaid on the $Z_{1}\text{-}Z_{3}$ panels.
@@
Figure 3 provides a numerical validation of the identifiability of the proposed UIDD. Specifically, Panel (b) demonstrates that the relative variation in the learned potential forces across different random seeds is significantly smaller for our method compared to standard OnsagerNet. Panels (c-h) further illustrate this by showing that UIDD consistently produces identical projected potential surfaces at $\dot{\varepsilon}=4.63$, irrespective of random initialization.
@@
Having established the identifiability, we now investigate the energy landscape of the polymer stretching dynamics under a range of strain rates $\dot{\varepsilon}$. The barrier height provides a quantitative measure of the energetic difficulty associated with transitions between metastable configurations [hanggi1990reaction, van1992stochastic]. In the context of polymer stretching, it represents the effective free-energy barrier associated with the flipping of chain ends against the imposed elongational flow. Although it does not exactly correspond to the transition rates in our nonequilibrium setting, it serves as a qualitative descriptor of the effective energy landscape. Since UIDD directly provides the logarithm of the stationary density, it allows for a straightforward computation of the barrier height (Methods D.2). As shown in Fig. 3 (i), the barrier height associated with polymer stretching exhibits a super-linear increase with the strain rate $\dot{\varepsilon}$. In contrast, the OnsagerNet approach fails to produce consistent energy functions for fixed parameters, resulting in barrier heights that show no clear or systematic dependence on the strain rate, as illustrated in Fig. 3 (j).
@@
Next, we examine the irreversibility by computing the global EPR, $\dot{S}_{\text{tot}}$, a key marker of non-equilibrium. The log-log plot in Fig. 4 (a) reveals a clear scaling relationship with the strain rate $\dot{\varepsilon}$. Specifically, $\dot{S}_{\text{tot}}$ exhibits sub-linear growth, tending towards a plateau in both the weak- and strong-driving regimes. The fact that $\dot{S}_{\text{tot}}$ is consistently positive for all strain rates provides evidence that the polymer system operates far from equilibrium.
@@
To understand the origin of this global irreversibility, we subsequently measure the total EPR. Fig. 4 (c-f) shows distinct dynamical regimes governed by the strain rate $\dot{\varepsilon}$ (see Methods D.4 for results at other $\dot{\varepsilon}$ values). At lower magnitudes of $\dot{\varepsilon}$, the total EPR exhibits minimal temporal variation along the individual trajectories of the polymers. In contrast, for larger values of $\dot{\varepsilon}$, the local EPR displays pronounced spatiotemporal heterogeneity during the unfolding process. Specifically, the total EPR is significantly suppressed at both the initial state (folded polymer) and the final state (fully extended conformation), yet it peaks dramatically during the mid-unfolding phase. This strain rate-dependent bimodal distribution of total EPR highlights the role of non-equilibrium dynamics in driven polymer unfolding, where maximum irreversibility occurs during mid-trajectory unfolding events.
@@
Finally, we conduct single-molecule real-world experiments to verify the qualitative results that the polymer system operates far from equilibrium. A property of the global EPR is that the EPR computed from projected dynamics is a lower bound for that of the full system (Methods D.1). Leveraging this property, we can assess non-equilibrium behavior directly from device-level observations by evaluating the EPR of the measured coordinates. We track the stretching dynamics of individual DNA molecules subjected to a planar elongational field, where the strain rate is regulated by the applied voltages (Methods D.4). Configurations in which the DNA chain is fully extended are selected to ensure that the system had reached a steady state. From the image-recorded configurations, we extract the time series of the horizontal coordinates of the left and right endpoints, corresponding to the atomic positions at both termini of the DNA molecule. Using only these two-dimensional trajectory data, we numerically compute the global EPR based on the definition (4). This procedure is illustrated in Fig. 4 (b). The resulting EPR values, $71.41$ at $60V$ and $68.99$ at $30V$, remain strictly positive even with partial information, consistent with the non-equilibrium behavior predicted by UIDD. These findings highlight the validity of our approach in extracting physically meaningful insights from real polymer dynamics.
@@ 2.4 Stochastic Gradient Langevin Dynamics
Stochastic gradient langevin dynamics (SGLD) [welling2011bayesian] is a cornerstone algorithm in modern machine learning, bridging stochastic optimization and Bayesian sampling. It extends standard mini-batch gradient descent by injecting isotropic Gaussian noise to sample from the target distribution:
@@
While the injected noise is designed to drive the system toward an equilibrium Gibbs-Boltzmann distribution, the stochasticity inherent in mini-batch gradient approximation complicates this picture. In the continuous-time limit, the mini-batching introduces a state-dependent covariance into the diffusion term [li2019stochastic]:
@@
We now apply our method to analyze SGLD in the context of least squares regression (linear) and independent components analysis (nonlinear) detailed in Methods D.5 to measure the irreversibility and validate the connection between irreversibility and sampling bias. Fig. 5 quantifies the impact of mini-batch size on SGLD irreversibility, measured by the global EPR, across varying batch sizes. Across two distinct systems, the global EPR consistently decreases with increasing mini-batch size at small scales, and plateaus at larger scales due to numerical precision constraints. These results indicate that increasing mini-batch size progressively suppresses non-equilibrium dynamics in SGLD. At sufficiently large batch sizes, negligible global total EPR values indicate that the SGLD performs near detailed balance, and that the non-equilibrium bias vanishes.
@@
Furthermore, we compute the MMD between the invariant distributions from different batch sizes and that of full-batch SGLD. As shown in Fig. 5 (a, c), the MMD trend as a function of batch size is highly consistent with that of the EPR (see Methods D.5 for theoretical justification). For comparison, we test a heuristic convergence metric based on the MMD between distributions from consecutive mini-batch sizes (i.e., batch sizes $b$ and $b/2$). This baseline metric can be misleading, showing trends inconsistent with actual convergence or lagging as an indicator, as shown in Fig. 5 (b, d). These findings demonstrate that UIDD constructed using only sampling trajectories can effectively compute the EPR as a physically meaningful diagnostic of sampling bias.
3 Discussion
@@
This work introduces UIDD, a framework that unifies physical interpretability, universal approximation, and rigorous identifiability for learning complex dissipative dynamics. It ensures a unique energy landscape and a clear decomposition of the drift force into reversible and irreversible components. This structure enables key thermodynamic quantities, particularly the entropy production rate that characterizes time-irreversibility, to be directly inferred from data. Applications to polymer stretching and stochastic gradient Langevin dynamics demonstrate the versatility of UIDD as a general tool for uncovering unknown dynamics in non-equilibrium phenomena. With its identifiable and physically consistent representation of dynamics, the framework opens new avenues for data-driven discovery and control of complex systems, offering a unified perspective on diverse nonequilibrium processes across scientific disciplines [dobson2003protein, whitesides2002self, zhao2023learning]. Building on these foundations, several promising directions remain for further exploration. A natural extension is to apply our approach to learn the non-equilibrium dynamics of other scientifically important systems, such as information processing in the brain [nartallo2026nonequilibrium], the thermodynamics of active matter [marchetti2013hydrodynamics], and protein folding [dobson2003protein]. Another direction is to design equilibrium-preserving sampler and numerical difference scheme by systematically quantifying and reducing irreversibility. Finally, our current framework inherently focuses on macroscopic dynamics learned from coarse-grained descriptions. Extending it to learn directly from mesoscopic or microscopic information through an operator-level formulation offers an exciting pathway toward broader and more versatile applications.
Data and Code Availability
The simulation and experimental datasets produced and analyzed for Polymer Stretching Dynamics are publicly available at the following Hugging Face repositories: https://huggingface.co/datasets/MLDS-NUS/Fs_T_Reduced, https://huggingface.co/datasets/MLDS-NUS/Fs_Validation_Reduced, and https://huggingface.co/datasets/MLDS-NUS/Experimental_Images, corresponding respectively to the primary simulation trajectories, validation simulation data, and validation experimental images. Further details on data generation are provided in Methods D.
The code used to reproduce the results is publicly available on GitHub at https://github.com/MLDS-NUS/UIDD-jax.
Acknowledgements
[…]
Methods
Appendix A Comparison to Existing Dissipative Dynamics Models
We compare UIDD with several thermodynamically motivated frameworks for learning dissipative dynamics.
Stat-PINNs
When compared with Stat-PINNs [huang_al_2025], UIDD can be viewed as their non-equilibrium extension. Stat-PINNs correspond precisely to the time-reversible (symmetric) sector of UIDD, recovering only the gradient-driven component of the dynamics. By additionally learning the antisymmetric operator $W$ and its divergence, UIDD captures the full irreversible drift and can therefore represent general non-equilibrium behavior. We note that directly adding a general antisymmetric operator $W$ (e.g. in OnsagerNet [yu2021onsagernet]) leads to non-identifiability, hence a careful construction of the banded structure of $W$ in UIDD is required.
S-OnsagerNet
Compared with S-OnsagerNet [chen2024constructing], which also incorporates non-equilibrium behavior, UIDD employs specific matrix functions for both $M$ and $W$ and includes their divergences as separate additive terms. These structures are not ad-hoc modifications but arise from a principled synthesis of the generalized Onsager principle [yu2021onsagernet], the generalized Helmholtz decomposition [da2023entropy], and the essentially Hamiltonian decomposition [feng1995volume], see Methods B.3. The resulting dynamical model provides an identifiable extension of S-OnsagerNet that, in addition, allows computation of both the invariant distribution and the EPR, as demonstrated in the applications.
Deterministic Models
UIDD naturally incorporates finite-temperature conditions, where temperature controls the noise intensity and influences thermodynamic quantities. To make this explicit, we introduce a parameter $T$ representing temperature and rescale
$\sigma=\sqrt{T}\,\tilde{\sigma},\qquad M=T\tilde{M}$,
together with the re-normalizations
$V=\tilde{V}/T,\qquad W=T\tilde{W}$.
Substituting these rescaled quantities into the original formulation yields the $T$-dependent form
Note that $\tilde{W}$ and $\tilde{V}$ still depend on $T$. Taking the limit $T\to 0$ and assuming that $\tilde{W}$ and $\tilde{V}$ have well-defined limits, UIDD formally reduces to the deterministic OnsagerNet dynamics [yu2021onsagernet],
$\dot{\mathbf{Z}}_{t}=-[M(\mathbf{Z}_{t})+W(\mathbf{Z}_{t})]\nabla V(\mathbf{Z}_{t}),$
as both the divergence terms and the noise vanish. In this limit, the formulation also coincides with the GENERIC formalism [lee2021machine, zhang2022gfinns],
$\dot{\mathbf{Z}}_{t}=L(\mathbf{Z}_{t})\,\nabla E(\mathbf{Z}_{t})+M(\mathbf{Z}_{t})\,\nabla S(\mathbf{Z}_{t}),\quad L(\mathbf{Z}_{t})\,\nabla V(\mathbf{Z}_{t})=M(\mathbf{Z}_{t})\,\nabla E(\mathbf{Z}_{t})=0,$
which can be written in Onsager form by taking V=(E+S)V=-(E+S) and $W=L$. It is noted that these deterministic models, and even models that merely add noise directly, are all non-identifiable. The key ingredients of UIDD, including the noise component that fixes $M$, the banded structure imposed on $W$, and the added divergence term, are all all essential for identifiability.
Classical Dissipative Formulations
Beyond neural-network-based approaches, there is a long-standing literature on thermodynamically consistent formulations of dissipative dynamics. Classical examples include Langevin equations [ayala2025reversibility], the Onsager principle [doi2011onsager, onsager1931reciprocal], the GENERIC formalism, and related potential-based formulations [ao2004potential]. These classical models specify the dissipative and conservative components of the dynamics in terms of prescribed energy, entropy, and transport operators. UIDD is designed to be compatible with this viewpoint: for fixed parameterizations of $V$, $M$, and $W$, it reduces to a standard SDE with well-defined thermodynamic structure, while its neural representation makes these objects learnable directly from data.
Overall Comparison
In summary, Stat-PINNs [huang_al_2025] correspond to the equilibrium, time-reversible sector with identifiability; S-OnsagerNet [chen2024constructing] extends this picture to stochastic dynamics without enforcing identifiability; and OnsagerNet or GENERIC based models [yu2021onsagernet, lee2021machine, zhang2022gfinns] arise as deterministic, zero-temperature limits. UIDD simultaneously extends these approaches to finite-temperature, stochastic, and non-equilibrium settings, yielding a mathematically universal and identifiable framework.
Appendix B Theoretical Results